A women is 5 times older than her son. In 3 years, she will be 4 times older than her son. How old is the women and her son?
Answer by Alan3354(69443)

A women is 5 times older than her son. In 3 years, she will be 4 times older than her son. How old is the women and her son?
============
5 times older = 6 times as old.
---
1 time older = 2x as old.
2 times older = 3x as old.
3 times older = 4x as old.
etc.
=============
W = 6S
W + 3 = 5(S + 3)
---
Sub for W
6S + 3 = 5S + 15
S = 12
W = 60